OpenStudy (anonymous):

I don't get this at all :(, here's the problem: 3AB + 2BA = B99 If A and B represent distinct digits correctly worked in the addition problem above, what is the value of A?

OpenStudy (jamesj):

You know from the 10 and 1 place that A + B = 9. And from the 100 place you have 3 + 2 = B, i.e., B = 5. Hence A = ...?

OpenStudy (anonymous):

um..i still do't get it o:

OpenStudy (anonymous):

3ab + 2ba = 99b 3a + 2a = 99 5a = 99 a = 99/5

OpenStudy (anonymous):

Haha I think I've misunderstood the question ENTIRELY.

OpenStudy (jamesj):

What we can do is look at each place: the units, the tens and the hundreds, and each of them gives us an equation. LHS = left-hand side = 3AB + 2BA = (3 * 100 + A * 10 + B * 1) + (2*100 + B*10 + A*1) = (2+3)*100 + (A+B)*10 + (B+A)*1 = 5*100 + (A+B)*10 + (B+A)*1 So far so good?

OpenStudy (anonymous):

I really have no idea D: where is the 399 from? Is this answer 153?

OpenStudy (jamesj):

RHS = B99 = B*100 + 9*10 + 9*1 Yes?

OpenStudy (jamesj):

right-hand side, is B99, not 399, and B99 is equal to B99 = B*100 + 9*10 + 9*1 yes?

OpenStudy (jamesj):

Hence the LHS = RHS if 5*100 + (A+B)*10 + (B+A)*1 = B*100 + 9*10 + 9*1 Yes?

OpenStudy (anonymous):

I'm extremely lost, james D:

OpenStudy (aravindg):

JAMES I NEED UR HELP BADLY

OpenStudy (aravindg):

PLZ COME AFTER THIS

OpenStudy (jamesj):

Ok. What does 123 mean? 123 = 1*100 + 2*10 + 3*1 do you agree?

OpenStudy (anonymous):

Since 3AB + 2AB = B99, we can deduce that B has to be 5. The limit of 3AB is 399, the limit of 2BA is 299 - the sum of which is 698, therefore it cannot be 6 as we know it ends in 99. So B = 5. From there, since 3+2 = 5, we can say that AB + BA = 99. Therefore find two two-digits numbers arranged in such a way that one is the reverse of the other, B represents 5 and they sum to 99 - 45, and 54 are the only matching pair. Therefore A = 4.

OpenStudy (anonymous):

but when you plug in those values, it doesn't work?

OpenStudy (jamesj):

it does work. B = 5, A = 4 works: 599 = 345 + 254

OpenStudy (anonymous):

OHHH they represent DIGITS IN the number. I was multiplying them.... /fail

OpenStudy (anonymous):

Sorry JamesJ I didn't mean to steal your thunder, you were doing an excellent job explaining it.

OpenStudy (jamesj):

no problem

OpenStudy (aravindg):

CAN U HLP ME JAMES?

OpenStudy (jamesj):

I left you a note on your problem.